Lavage Rapide is a Canadian company that owns and operates a large automatic car wash facility near Montreal. The following table provides data concerning the company’s costs:

Fixed Cost
per Month
Cost per
Car Washed
Cleaning supplies $ 0.80
Electricity $ 1,200 $ 0.15
Maintenance $ 0.20
Wages and salaries $ 5,000 $ 0.30
Depreciation $ 6,000
Rent $ 8,000
Administrative expenses $ 4,000 $ 0.10

For example, electricity costs are $1,200 per month plus $0.15 per car washed. The company expects to wash 9,000 cars in August and to collect an average of $4.90 per car washed.
Required:

Prepare the company’s planning budget for August.
Answer #1
Planning Budget
Budgeted Cars washed 9000
Revenue 44100 =9000*4.9
Expenses:
Cleaning supplies 7200 =9000*0.8
Electricity 2550 =1200+(9000*0.15)
Maintenance 1800 =9000*0.2
Wages and salaries 7700 =5000+(9000*0.3)
Depreciation 6000
Rent 8000
Administrative expenses 4900 =4000+(9000*0.1)
Total expenses 38150
Net operating income 5950